OpenStudy (anonymous):

the measure of the supplemt of an angle is 40degree less than 6 times the measure of its complement . Find the angle . ( use algebra) How to do this problem ?? Show steps please !! >< All answer gets medals !!

OpenStudy (ash2326):

Let the angle be x We are given the supplement of the angle (180-x) is 40 less than six times its complement (90-x) so we have \[(180-x) =6(90-x)-40\] so we get now \[180-x=540-6x-40\] Can you solve now????

OpenStudy (ash2326):

@MetalPencil did you understand? Can you do it now?

OpenStudy (anonymous):

yup !!! thanks a LOT :)

OpenStudy (ash2326):

Welcome:D